OpenStudy (anonymous):

Given a system of three linear equations in three variables where y = 8 and D = −4 (D is the determinant of the coefficient matrix), use Cramer’s Rule to find Dy.

OpenStudy (anonymous):

-32 ? o.O

OpenStudy (anonymous):

huh how u get the one before and how sure are u on that one?

OpenStudy (anonymous):

Cramer's rule for three variables of x,y,z would be: x=Dx/D; y=Dy/D; z=Dz/D; the above answer of -32 is correct.